# Neural Audio Codecs for Prompt-Driven Universal Source Separation

### v1

@@ 5 Conclusion
Tables 8(a)8(d) consolidate relative improvements of CodecSep over AudioSep under matched training data and prompt protocols, complementing the absolute results in the main text. On DnR-v2 (Table 8(a)), CodecSep delivers large SI-SDR gains—especially for music and SFX—together with a clear perceptual lift. Under paraphrased prompts (Table 8(b)), gains are smaller but remain positive, indicating robustness to lexical variation. Across additional open-domain benchmarks (Table 8(c)), SI-SDR gains are consistent while ViSQOL deltas are modest, aligning with cross-domain trends reported earlier. Finally, when trained on AudioCaps (Table 8(d)), CodecSep maintains an advantage on AudioCaps-test and yields strong improvements on DnR-v2, supporting the claim that codec-latent masking generalizes well across datasets and prompt regimes.
@@
Appendix B Further Studies: Reconstruction Performance.

### v2

@@ Appendix C Training Details
Tables 8(a)8(d) consolidate relative improvements of CodecSep over AudioSep under matched training data and prompt protocols, complementing the absolute results in the main text. On dnr-v2 (cf. Table 8(a)), CodecSep delivers large SI-SDR gains—especially for music and SFX—together with a clear perceptual lift. Under paraphrased prompts (cf. Table 8(b)), gains are smaller but remain positive, indicating robustness to lexical variation. Across additional open-domain benchmarks (cf. Table 8(c)), SI-SDR gains are consistent while ViSQOL deltas are modest, aligning with cross-domain trends reported earlier. Finally, when trained on AudioCaps (cf. Table 8(d)), CodecSep maintains an advantage on AudioCaps-test and yields strong improvements on dnr-v2, supporting the claim that codec-latent masking generalizes well across datasets and prompt regimes.
@@
Appendix F Further Studies: Reconstruction Performance.